# Iterative Trajectory Exploration for Multimodal Agents

### v1

@@ Abstract
Multimodal agents, which integrate a controller (e.g., a large language model) with external tools, have demonstrated remarkable capabilities in tackling complex tasks. However, existing agents need to collect a large number of expert data for fine-tuning to adapt to new environments. In this paper, we propose an online self-exploration method for multimodal agents, namely SPORT, via step-wise preference optimization to refine the trajectories of agents, which automatically generates tasks and learns from solving the generated tasks, without any expert annotation. SPORT operates through four iterative components: task synthesis, step sampling, step verification, and preference tuning. First, we synthesize multimodal tasks using language models. Then, we introduce a novel search scheme, where step sampling and step verification are executed alternately to solve each generated task. We employ a verifier to provide AI feedback to construct step-wise preference data. The data is subsequently used to update the controller’s policy through preference tuning, producing a SPORT Agent. By interacting with real environments, the SPORT Agent evolves into a more refined and capable system. Evaluation in the GTA and GAIA benchmarks shows that the SPORT Agent achieves 6.41% and 3.64% improvements, underscoring the generalization and effectiveness introduced by our method. The project page is https://SPORT-Agents.github.io.
Machine Learning, ICML
@@ 1 Introduction
Leveraging large language models (LLMs) or vision-language models (VLMs) as controllers to call external tools has become a promising direction in building multimodal agents (Surís et al., 2023; Gupta & Kembhavi, 2023; Zhong et al., 2023; Wang et al., 2024a), achieving impressive performance across multiple complex downstream tasks, such as GUI control (Cheng et al., 2024), multimodal reasoning (Gao et al., 2024a), and embodied AI (Fan et al., 2024). To enhance the planning and reasoning abilities of agents in new environments, existing studies focus on collecting training data to tune the controller of an agent, where commonly used techniques include supervised fine-tuning (SFT) (Hu et al., 2024; Gao et al., 2024b) and reinforcement learning (RL) (Deng et al., 2024c; Xiong et al., 2024b) for alignments with desired behaviors.
The key to agent tuning is collecting sufficient expert data (e.g., tasks, labels, and trajectories to solve tasks). SFT methods collect expert data via human annotation or distillation from closed-source API to fine-tune the controllers. RL methods use the expert data to tune reward models, update policy models (Qi et al., 2024; He et al., 2024), or derive performance data for optimization (Zhang et al., 2024; Putta et al., 2024). However, collecting high-quality expert data may be difficult in new environments. It is labor-extensive and high-cost, and such pre-collected data may lead to biased distributions inconsistent with the target environments, causing inferior results.
@@
In this paper, we focus on the multimodal reasoning tasks, where the agents are required to call diverse tools (e.g., web search, visual reasoning, file understanding, and object localization) to answer given questions. We explore whether multimodal agents can improve their performance via online self-exploration without any expert data. We are inspired by existing research in LLMs and VLMs, which has shown impressive performance in self-instruction (Wang et al., 2023; Liu et al., 2023), self-verification (Madaan et al., 2024; Yu et al., 2024), and self-learning (Deng et al., 2024b; Kumar et al., 2024). Based on the above observation, we expect that the agent automatically generates tasks, searches for possible trajectories to call tools in solving synthetic tasks, evaluates these trajectories by itself, and updates controllers using this data. In this case, agents will improve their generalization capability by interacting with environments.
To achieve this goal, we must address three key challenges in such an online self-exploration framework for multimodal reasoning tasks. (1) Lack of off-the-shelf tasks and expert trajectories. There is no off-the-shelf dataset with ground truth annotation for multimodal reasoning tasks. It is non-trivial to identify whether a task is correctly solved, thus making it challenging to select correct trajectories. (2) Difficulty in reward modeling. Unlike other settings (e.g., GUI setting), where sufficient data or pre-defined rules make reward modeling feasible, predicted answers or trajectories in multimodal reasoning cannot be easily and reliably verified. (3) Low sampling efficiency and high cost. Sampling trajectories often involves expensive tools (e.g., LLM APIs, web search), which results in both high monetary and computational costs, making it challenging to scale up.
@@
To solve the above challenges, we propose SPORT, an iterative self-exploration method via step-wise preference optimization to refine trajectories of multimodal agents, as shown in Figure 1. SPORT operates through four iterative components: task synthesis, step sampling, step verification, and preference tuning. First, we generate queries and multimodal files for task synthesis based on provided task seeds. Second, we introduce a new search scheme that samples step-level candidate solutions to call tools. Third, we employ a multimodal verifier that, given the task context, intermediate reasoning states, and step candidates, provides AI-generated feedback to estimate step-level preferences. Finally, we perform step-wise preference tuning to refine the controller’s policy and obtain the SPORT Agent, which is then used to guide trajectory sampling in the next iteration.
@@
SPORT enables agents to autonomously generate tasks and explore tool-use trajectories, removing reliance on expert data or pre-collected datasets. For pre-trained LLMs, providing step-level preference signals is easier than accurate rewards, circumventing the difficulties of reward modeling in complex multimodal tasks. Furthermore, SPORT improves the utilization of sampled trajectories—even for some failed trajectories, we can still discover useful step-level preference data, allowing us to acquire more effective data utilization with the same number of sampled trajectories. These capabilities collectively support stable, scalable self-exploration for complex multimodal tasks.
@@
In summary, our contributions are three-fold.
(1) We propose an online self-exploration framework for multimodal agents, through which the agents can adapt to new environments without any expert annotation.
(2) We propose SPORT, an online self-exploration method that utilizes step-wise optimization and AI feedback, providing a possible way to self-learning in complex multimodal environments.
(3) The obtained SPORT Agent achieves significant performance improvements compared with SFT based agents on two popular benchmarks: GTA and GAIA.
@@ 2.1 Agent Tuning
Due to the disparity between the LLMs and the requirements of agents, agent tuning is necessary to adapt to practical tasks in new environments. Research for agent tuning could be divided into two categories: supervised fine-tuning (SFT) and reinforcement learning (RL). SFT methods collect expert trajectory data via distillation from closed-source API (e.g., GPT-4o) (Gao et al., 2024b; Liu et al., 2024d; Zeng et al., 2023) or human annotation (Liu et al., 2024c; Deng et al., 2024a). Then they use these collected data to tune the controller via SFT. However, the SFT methods suffer from huge costs and inferior generalization (Shi et al., 2024). To solve this issue, researchers have paid attention to RL agent tuning methods that allow agents to interact with the environment and learn from the feedback. Some methods utilize the policy gradient technique (Zhou et al., 2024a) to update the controller with a reward model that is designed as a fine-tuned model (Qi et al., 2024; Zhai et al., 2024b), environment feedback (Bai et al., 2024; Zhai et al., 2024a), human-designed rules (He et al., 2024; Zhou et al., 2024b), or tree search results (Deng et al., 2024c). To simplify this procedure, the direct preference optimization (DPO) method is applied to agent tuning (Xiong et al., 2024b), where trajectory-level (Zhang et al., 2024; Song et al., 2024) or step-level preference data (Putta et al., 2024; Chen et al., 2024b) is constructed to update the controller. The above DPO methods construct step-wise preference data based on whole correct trajectories. Nevertheless, such expert or correct trajectories are difficult to obtain in complex multimodal tasks, and the predefined correct trajectories may lead to biased distributions. In contrast, our online self-exploration framework does not rely on any correct trajectories. In addition, it uses an online manner to avoid biased distributions.
@@ 2.2 Step-wise Preference Tuning
Preference tuning methods rely on paired data, which is not readily available for complex tasks with multi-step reasoning, making it non-trivial to determine which trajectory is better. Furthermore, for long trajectories, only an overall preference verification can not capture the relationships among steps and ignores the fine-grained preference between different steps. To overcome this problem, step-wise preference has been studied. STEP-DPO (Lai et al., 2024) and SCDPO (Lu et al., 2024) collect step-wise preference data by localizing error steps or disturbing the correct path. OREO (Chen et al., 2024a) and SVPO (Wang et al., 2024b) train value models for step-wize verification and inference guidance. SDPO (Kong et al., 2025) combines step-level, turn-level, and session-level preference data for full-grained optimization. The above methods mainly focus on code generation and math reasoning tasks that are easy to obtain correct trajectories to construct step-wise preference data. In contrast, this method focuses on multimodal agents, where obtaining correct trajectories is challenging. Thus, our agent explores the environment by itself via an online manner, which uses designed AI feedback to construct preference data without any annotation.
@@ 2.3 Learning from AI Feedback
Using models to generate AI feedback (natural languages, scores, or preference) for performance improvement has emerged as a critical paradigm (Bai et al., 2022). Existing methods can be broadly divided into three categories based on how to use the obtained AI feedback. Methods of the first category add the AI feedback into prompts for in-context learning (Madaan et al., 2024). LLaVA-Cirtic (Xiong et al., 2024a) collects sufficient data to train a model to provide rich multimodal AI feedback. VLM-F (Liao et al., 2024), VolCaNo (Lee et al., 2024a), and Clarify (Lee et al., 2024b) add AI feedback to VLM prompts to address visual hallucinations. CLOVA (Gao et al., 2024a) and CompAgent (Wang et al., 2024f) are representative agents that use AI feedback to refine prompts for better task-solving. Methods of the second category use AI feedback to filter data for supervised fine-tuning. M-STAR (Liu et al., 2024b) shows its effectiveness in visual mathematical reasoning. APIGen (Liu et al., 2024d), MAT (Gao et al., 2024b), and visualagentbench (Liu et al., 2024c) design verifiers to collect trajectories for agents. Methods of the third category use AI feedback for reinforcement learning. They use LLMs or VLMs to produce rewards for policy gradient optimization (Lee et al., ; Fu et al., ; Rocamonde et al., 2023), and several of them produce preference data for the DPO algorithm (Yu et al., 2024). Different from them, our AI feedback is well-designed for evaluating each step in complex scenarios of multimodal agents.
Figure 2: Demonstrations of the search scheme used in the SPORT method. Given a task, the agent samples potential solutions for each step and verifies their qualities in an online manner. Then, we construct the step-wise preference data based on such self-exploration.
@@ 3.1 Formulation
We opt for the framework of the ReAct agent (Yao et al., 2023) that performs step-by-step reasoning for tool usage for task solving. In each step, based on the input $x_{i}$, the agent outputs a solution $s_{i}$ for tool calling.
$s_{i}^{\star}=\arg\max\pi_{\theta}(s_{i}|x_{i},T),$ (1)
where $\pi_{\theta}$ is the controller (an VLM in our method) of agents with $\theta$ being the parameters, $x_{i}$ is composed of the task (including a query $Q$ in natural language and multimodal files $F$) and the history $h_{i}$ of previous steps, i.e., $x_{i}=\{Q,F,h_{i}\}$. The solution $s_{i}$ consists of the thought $t_{i}$ and code $c_{i}$, $s_{i}^{=}\{t_{i},c_{i}\}$. $T$ denotes available tools, and we follow the work (Gao et al., 2024b) using the same toolkit. In this case, we further rewrite Eq. (1) as
@@
where $t_{i}^{\star}$ and $c_{i}^{\star}$ are thought and code for the $i$-th step, and the history $h_{i}=\{t_{1},c_{1},o_{1},\cdots,t_{i-1},c_{i-1},o_{i-1}\}$ is composed of thought, code, and observation of previous steps.
Agent tuning aims to update $\theta$ to increase the planning and reasoning capabilities of agents in new environments. This paper proposes an online self-exploration method, SPORT, to update $\theta$ via step-wise preference optimization in refining trajectories, as shown in Figure 1. Concretely, SPORT has iterative components: task synthesis, step sampling, step verification, and preference tuning. In one iteration, SPORT first generates some multimodal tasks. For each generated task, SPORT performs step sampling and step verification alternately to construct step-wise preference data. Finally, SPORT uses the step-wise preference data to tune the controller.
@@ 3.2 Task Synthesis
Since a multimodal task is composed of a language query and multimodal files, the task synthesis component is divided into query generation and multimodal file generation. We first generate queries and then generate files, rather than the reverse order, since the multimodal files are diverse (such as DOCX, PPTX, XLSX, and PDF are commonly encountered), and it is challenging to construct a diverse file dataset in advance. In addition, tasks are usually based on multiple files instead of only one. First obtaining files and then generating queries may cause weak relevance of files and produce non-natural queries.
To produce diverse and practical queries, we collect seed from the existing method MAT (Gao et al., 2024b) and employ an LLM (GPT-4o-mini in practice) to generate queries. We feed randomly sampled seed queries, used tools, and a designed prompt to the LLM that generates multiple queries once. Adding tool descriptions to the prompt makes GPT-4o mini better understand what queries can be solved, improving their qualities. For each generated query, we prompt GPT-4o mini to output the needed files. If images are needed, we search for source images from off-the-shelf datasets based on similarities. For other files, we prompt GPT-4o mini to generate Python code to produce files.
@@ 3.3 Data construction
The two components: step sampling and step verification are performed to construct high-quality preference data, playing key roles in our method. To avoid potential bias issues in constructed data, we introduce an online search scheme, where the step sampling and step verification are executed alternately in each generated task, as shown in Figure 2.
The step-wise preference data is formulated as a triplet $(x_{i},s_{i}^{pre},s_{i}^{dis})$, where $x_{i}=\{Q,F,h_{i}\}$ denotes the input including the query $Q$, files $F$, and the history $h_{i}$ of previous steps, $s_{i}^{pre}$ is the preferred solution in the current step including, and $s_{i}^{dis}$ is the dispreferred solution.
Given a task with the query $Q$ and files $F$, the agent expands the search space for the first step by sampling $n$ solutions $\{s_{1}^{1},s_{1}^{2},\cdots,s_{1}^{n}\}$, including the thought and code $\{t_{1}^{1},c_{1}^{1},\cdots,t_{1}^{n},c_{1}^{n}\}$ from the controller, and execute them to obtain $n$ observations $\{o_{1}^{1},\cdots,o_{1}^{n}\}$. Then, we feed the query $Q$, $n$ solutions, and $n$ observations to an LLM, and ask the LLM to select the best solution $\{t_{1}^{*},c_{1}^{*}\}$ with its corresponding observation $o_{1}^{*}$.
Along $\{t_{1}^{*},c_{1}^{*}\}$ and $o_{1}^{*}$, we expand the search space for the second step. Regarding $\{t_{1}^{*},c_{1}^{*},o_{1}^{*}\}$ as the history $h_{2}$, the controller samples $n$ solutions $\{t_{2}^{1},c_{2}^{1},\cdots,t_{2}^{n},c_{2}^{n}\}$ from the controller, and executes them to obtain $n$ observations $\{o_{2}^{1},\cdots,o_{2}^{n}\}$. We feed the query $Q$, history $h_{2}$, $n$ solutions in this step, and $n$ corresponding observations to an LLM, and ask the LLM to select the best solution $\{t_{2}^{*},c_{2}^{*}\}$ with its corresponding observation $o_{2}^{*}$. In this case, the agent gradually expands the search space and selects the best solution for each step, until the agent believes that the task is over.
Assume there are $m$ steps in solving one task. In this case, we could collect $m(n-1)$ preference data pairs. In the $i$-th step, the selected best solution $\{t_{i}^{*},c_{i}^{*}\}$ is the preferred output, and the rest $n-1$ solutions are the dispreferred outputs, denoted as $\mathcal{D}_{i}^{dis}$, $|\mathcal{D}_{i}^{dis}|=n-1$. The preference data in one task is denoted as $\{(x,\{t_{i}^{*},c_{i}^{*}\},\{t_{i}^{j},c_{i}^{j}\})\}$, where $\{t_{i}^{j},c_{i}^{j}\}\in\mathcal{D}_{i}^{dis}$ and $i\in[1,m]$.

### v4

# Iterative Tool Usage Exploration for Multimodal Agents via Step-wise Preference Tuning

@@ Abstract
Multimodal agents, which integrate a controller (e.g., a vision language model) with external tools, have demonstrated remarkable capabilities in tackling complex multimodal tasks. Existing approaches for training these agents, both supervised fine-tuning and reinforcement learning, depend on extensive human-annotated task-answer pairs and tool trajectories. However, for complex multimodal tasks, such annotations are prohibitively expensive or impractical to obtain. In this paper, we propose an iterative tool usage exploration method for multimodal agents without any pre-collected data, namely SPORT, via step-wise preference optimization to refine the trajectories of tool usage. Our method enables multimodal agents to autonomously discover effective tool usage strategies through self-exploration and optimization, eliminating the bottleneck of human annotation. SPORT has four iterative components: task synthesis, step sampling, step verification, and preference tuning. We first synthesize multimodal tasks using language models. Then, we introduce a novel trajectory exploration scheme, where step sampling and step verification are executed alternately to solve synthesized tasks. In step sampling, the agent tries different tools and obtains corresponding results. In step verification, we employ a verifier to provide AI feedback to construct step-wise preference data. The data is subsequently used to update the controller for tool usage through preference tuning, producing a SPORT agent. By interacting with real environments, the SPORT agent gradually evolves into a more refined and capable system. Evaluation in the GTA and GAIA benchmarks shows that the SPORT agent achieves 6.41% and 3.64% improvements, underscoring the generalization and effectiveness introduced by our method.
@@ 1 Introduction
Leveraging large language models (LLMs) or vision-language models (VLMs) as controllers to call external tools (e.g., web search, visual reasoning, file understanding, and object localization) has become a promising direction in building multimodal agents (suris2023vipergpt, ; gupta2023visual, ; fan2024videoagent, ; wang2024tool, ), achieving impressive performance for complex tasks (gao2024clova, ; liu2025visual, ; li2025mm, ; fan2025eva, ). To enhance the planning and reasoning abilities of agents, existing studies focus on collecting tool usage trajectories to fine-tune the controller of an agent (hu2024visual, ; gao2024multimodalagenttuningbuilding, ), using human annotation or distillation from closed-source APIs. However, collecting high-quality tool usage data is labor-intensive and high-cost, and such pre-collected data may lead to biased distributions inconsistent with the target environments (such as task distributions and available tools), causing inferior generalization.
@@
In this paper, we study whether multimodal agents can improve their tool usage capability via self-exploration without any pre-collection. We are inspired by existing research in LLMs and VLMs, which has shown impressive performance in self-instruction wang2023self ; liu2024visual , self-verification madaan2023self ; yu2024rlaif , and self-learning deng2024enhancing ; kumar2024training . Based on the above observation, we expect that agents automatically generates tasks, searches for useful tools in solving synthetic tasks, evaluates the exploration by itself, and updates controllers using the exploration process. In this case, agents will improve the generalization of tool usage by interacting with environments.
To achieve this goal, we must address two key challenges in such tool usage exploration for complex multimodal tasks. (1) Lack of off-the-shelf tasks and annotation. There is no off‑the‑shelf dataset with ground truth annotations (answers and trajectories) for solving multimodal tasks. These tasks are usually domain-specific wang2024gta ; mialon2023gaia , making it non-trivial to verify whether the outcome is correct. Solving these tasks requires a long trajectory to call diverse tools, and thus, it is challenging to produce an exactly correct trajectory and compare the quality among different trajectories. (2) Low sampling efficiency and high cost. The exploration process requires executing the sampled tools (e.g., large language models, web search, and image generation), which results in both high monetary and computational costs, making it challenging to scale up.
@@
To solve the above challenges, we propose SPORT, an iterative tool usage exploration method via step-wise preference optimization to refine trajectories of multimodal agents, as shown in Figure˜1. SPORT operates through four iterative components: task synthesis, step sampling, step verification, and preference tuning. Firstly, we generate queries and multimodal files for task synthesis based on provided task seeds. Secondly, we introduce a new search scheme that samples step-level candidate actions (including the thoughts and codes) to call tools. Thirdly, we employ a multimodal verifier that, given the task context, step actions with execution results, provides AI-generated feedback to estimate step-level preferences. Finally, we perform step-wise preference tuning to refine the controller and obtain the SPORT agent, which is then used to guide tool sampling in the next iteration.
@@
SPORT enables agents to autonomously generate tasks and explore tool usage trajectories, removing reliance on pre-collected datasets. Step-level verification is easier than trajectory-level evaluation for pre-trained LLMs, circumventing annotation difficulties. Furthermore, SPORT improves data utilization by extracting useful step-level preferences even from failed trajectories, enabling more effective learning with the same number of samples. These capabilities enable stable and scalable self-exploration, achieving better generalization for complex multimodal tool usage tasks.
@@
In summary, our contributions are threefold. (1) We propose SPORT, a tool usage exploration framework for multimodal agents, providing a possible way for multimodal tool learning, leading to generalization without any annotation. (2) The obtained SPORT Agent achieves significant performance improvements compared with SFT agents on two popular benchmarks: GTA and GAIA. (3) We collect the explored preference tuning data into a dataset composed of 16K data, which is conducive to the subsequent research on tool usage and multimodal agents.
@@ 2.1 Agent Tuning
Due to the disparity between the LLMs and the requirements of agents, agent tuning is necessary to adapt to practical tasks. Research for agent tuning could be divided into two categories: supervised fine-tuning (SFT) and reinforcement learning (RL). SFT methods collect trajectory data via distillation from closed-source API (e.g., GPT-4o) gao2024multimodalagenttuningbuilding ; liu2024apigen ; zeng2024agenttuning ; zhang2025tongui or human annotation liu2024visualagentbench ; deng2024mimir . Then they use these collected data to tune the controller via SFT. However, the SFT methods suffer from huge costs and inferior generalization shi2024direct . To solve this issue, researchers have paid attention to RL agent tuning methods that allow agents to interact with the environment and learn from the feedback. Some methods utilize the policy gradient technique zhou2024proposer to update the controller with a reward model that is designed as a fine-tuned model qi2024webrl ; zhai2025enhancing , environment feedback bai2024digirl ; zhai2024fine , human-designed rules peiyuan2024agile ; zhou2024archer , or tree search results deng2024novice . Especially, some methods resort to the policy gradient method for tool learning jin2025search ; feng2025retool , where they use the prediction correctness as the reward. To simplify this procedure, the direct preference optimization (DPO) methods are applied to agent tuning xiong2024watch , which construct step-level putta2024agent ; chen2024advancing or trajectory-level zhang2024grape ; song2024trial preference data based on whole correct trajectories. Nevertheless, most existing agent tuning methods rely on answer or trajectory annotations that are difficult to obtain in multimodal tool usage tasks. In contrast, our tool usage exploration framework does not rely on any annotation via step-wise preference tuning.
@@ 2.2 Step-wise Preference Tuning
Preference tuning methods rely on paired data, which is not readily available for complex tasks with multi-step reasoning, making it non-trivial to determine which trajectory is better. Furthermore, for long trajectories, only an overall preference verification can not capture the relationships among steps and ignores the fine-grained preference between different steps. To overcome this problem, step-wise preference has been studied. STEP-DPO lai2024step and SCDPO lu2024step collect step-wise preference data by localizing error steps or disturbing the correct path. OREO chen2024step and SVPO wang2024offline train value models for step-wise verification and inference guidance. SDPO kong2025sdpo combines step-level, turn-level, and session-level preference data for full-grained optimization. The above methods mainly focus on code generation and math reasoning tasks that are easy to obtain correct trajectories to construct step-wise preference data. In contrast, this method focuses on multimodal agents for tool usage, where obtaining correct trajectories is challenging. Thus, our agent explores the tool usage by itself via an iterative manner, which uses designed AI feedback to construct step-wise preference data without any annotation.
@@ 2.3 Learning from AI Feedback
Using models to generate AI feedback for performance improvement has emerged as a critical paradigm bai2022constitutional . Existing methods can be broadly divided into three categories. The first category adds AI feedback into prompts for in-context learning madaan2023self . LLaVA-Cirtic xiong2024llava trains a model to provide multimodal AI feedback. VLM-F liao2024can , VolCaNo lee2024volcano , and Clarify lee2024clarify use AI feedback to address visual hallucinations. CLOVA gao2024clova and CompAgent wang2024divide refine prompts using AI feedback. The second category uses AI feedback to filter data for supervised fine-tuning, such as M-STAR liu2024diving for visual mathematical reasoning and APIGen liu2024apigen , MAT gao2024multimodalagenttuningbuilding , and visualagentbench liu2024visualagentbench for agent trajectories. The third category uses AI feedback for reinforcement learning, producing rewards for policy gradient optimization leerlaif ; fufurl ; rocamonde2023vision or preference data for DPO yu2024rlaif . Different from them, our AI feedback is well-designed for evaluating tool usage in complex scenarios of multimodal agents.
@@ 3 Method
Figure 2: Demonstrations of the search scheme used in the SPORT method. Given a task, the agent samples potential actions for each step and verifies their qualities in an online manner. Then, we construct the step-wise preference data based on such self-exploration.
@@ 3.1 Formulation
We opt for the framework of the ReAct agent yao2023react that performs step-by-step reasoning for tool usage. In each step, based on the input $x_{i}$, the agent outputs an action $a_{i}$ for tool calling.
$a_{i}^{\star}=\arg\max\pi_{\theta}(a_{i}|x_{i},T),$ (1)
where $\pi_{\theta}$ is the controller (an VLM in our method) of agents with $\theta$ being the parameters, $x_{i}$ is composed of the task (including a query $Q$ in natural language and multimodal files $F$) and the history $h_{i}$ of previous steps, i.e., $x_{i}=\{Q,F,h_{i}\}$. The action $a_{i}$ consists of the thought $t_{i}$ and code $c_{i}$ for tool calling, $a_{i}=\{t_{i},c_{i}\}$. $T$ denotes available tools, and we follow the work gao2024multimodalagenttuningbuilding using the same toolkit. In this case, we further rewrite Eq. (1) as
@@
where $t_{i}^{\star}$ and $c_{i}^{\star}$ are thought and code for the $i$-th step, and the history $h_{i}=\{t_{1},c_{1},o_{1},\cdots,t_{i-1},c_{i-1},o_{i-1}\}$ is composed of thought $\{t_{1,...,i-1}\}$, code $\{c_{1,...,i-1}\}$, and observation $\{o_{1,...,i-1}\}$ of previous steps.
Agent tuning aims to update $\theta$ to increase the tool usage capabilities of agents. This paper proposes an iterative tool usage exploration method, SPORT, to update $\theta$ via step-wise preference optimization in refining trajectories, as shown in Figure˜1. Concretely, SPORT has iterative components: task synthesis, step sampling, step verification, and preference tuning. In one iteration, SPORT first generates some multimodal tasks. For each generated task, SPORT performs step sampling and step verification alternately to construct step-wise preference data. Finally, SPORT uses the step-wise preference data to tune the controller.
@@ 3.2 Task Synthesis
Since a multimodal task is composed of a language query and multimodal files, the task synthesis component is divided into query generation and multimodal file generation. We first generate queries and then generate files, rather than the reverse order, since the multimodal files are diverse (such as DOCX, PPTX, XLSX, and PDF), and it is challenging to construct a diverse file dataset in advance. In addition, tasks are usually based on multiple files instead of only one. First obtaining files and then generating queries may cause weak relevance of files and unnatural queries.
To produce diverse and practical queries, we collect seed from the existing method MAT gao2024multimodalagenttuningbuilding and employ an LLM (e.g., Qwen2.5-7B) to generate queries. We feed randomly sampled seed queries, used tools, and a designed prompt to the LLM that generates multiple queries at once. For each generated query, we prompt the LLM to output the needed files. If images are needed, we search for source images from off-the-shelf datasets based on similarities. For other files, we prompt the LLM to generate Python code to produce files.
@@ 3.3 Data Construction
The two components: step sampling and step verification, are performed to construct high-quality preference data, playing key roles in our method. To avoid potential bias issues in constructed data, we introduce an online search scheme, where the step sampling and step verification are executed alternately in each generated task, as shown in Figure˜2.
The step-wise preference data is formulated as a triplet $(x_{i},a_{i}^{pre},a_{i}^{dis})$, where $x_{i}=\{Q,F,h_{i}\}$ denotes the input including the query $Q$, files $F$, and the history $h_{i}$ of previous steps, $a_{i}^{pre}$ is the preferred action in the current step, and $a_{i}^{dis}$ is the dispreferred action.
Concretely, given a task with the query $Q$ and files $F$, the agent expands the search space for the first step by sampling $n$ actions $\{a_{1}^{1},a_{1}^{2},\cdots,a_{1}^{n}\}$, including the thought and code $\{t_{1}^{1},c_{1}^{1},\cdots,t_{1}^{n},c_{1}^{n}\}$ from the controller, and execute them to obtain $n$ observations $\{o_{1}^{1},\cdots,o_{1}^{n}\}$. Then, we feed the query $Q$, $n$ actions, and $n$ observations to an LLM, and ask an LLM to select the best action $\{t_{1}^{*},c_{1}^{*}\}$ with its corresponding observation $o_{1}^{*}$.
Along $\{t_{1}^{*},c_{1}^{*}\}$ and $o_{1}^{*}$, we expand the search space for the second step. Regarding $\{t_{1}^{*},c_{1}^{*},o_{1}^{*}\}$ as the history $h_{2}$, the controller samples $n$ actions $\{t_{2}^{1},c_{2}^{1},\cdots,t_{2}^{n},c_{2}^{n}\}$ from the controller, and executes them to obtain $n$ observations $\{o_{2}^{1},\cdots,o_{2}^{n}\}$. We feed the query $Q$, history $h_{2}$, $n$ actions in this step, and $n$ corresponding observations to the LLM, and ask the LLM to select the best action $\{t_{2}^{*},c_{2}^{*}\}$ with its corresponding observation $o_{2}^{*}$. In this case, the agent gradually expands the search space and selects the best action for the next step, until the agent believes that the task is over. Here, we provide observations $\{o_{2}^{1},\cdots,o_{2}^{n}\}$ and prompt the verifier to check which tools lead to desirable observations, instead of verifying which observation is correct.
Assume there are $m$ steps in solving one task. In this case, we could collect m(n1)m(n-1) preference data pairs. In the $i$-th step, the selected best action $\{t_{i}^{*},c_{i}^{*}\}$ is the preferred output, and the rest n1n-1 actions are the dispreferred outputs, collected in a set $\mathcal{D}_{i}^{dis}$, $|\mathcal{D}_{i}^{dis}|=n-1$. The preference data in one task is denoted as $\{(x_{i},a_{i}^{pre},a_{i}^{dis})\}$, where $a_{i}^{pre}=\{t_{i}^{*},c_{i}^{*}\}$, $a_{i}^{dis}=\{t_{i}^{j},c_{i}^{j}\}\in\mathcal{D}_{i}^{dis}$, and $i\in[1,m]$.